OpenStudy (anonymous):

a car traveling at 7.0 m/s accelerates uni formally at 7.0 m/s accelerates uniformly at the rate of 0.80m/s^2 for a distance of 245 m. what is the velocity at the end of the acceleration?

OpenStudy (anonymous):

you can use equation \[v^2=v_0^2+2ax\]

OpenStudy (anonymous):

what is the v power 2 over 0?

OpenStudy (anonymous):

That's v sub zero squared, the initial velocity squared.

OpenStudy (anonymous):

so what is the rule i use?

OpenStudy (anonymous):

@raper18 , your question is unclear due to a few grammatical errors, but I believe \[v_0 = 7 m/s \\ a= .8 m/s^2 \\ x = 245 m\] Then using @bourne13 's equation, you can solve for v, the final velocity. Also, as a side note, if it's possible/while there isn't much associated with your account, you *very much* might want to add an extra p in your name - ie, rapper18 . As it stands, your account name is very unpleasant when read in English.

OpenStudy (anonymous):

Two good points.